# BoostDream: Efficient Refining for High-Quality Text-to-3D Generation from Multi-View Diffusion

### v3

@@ Abstract
Figure 1: Comparison of 3D Generation Results of baseline and BoostDream. Provided with a coarse 3D asset and text prompt pair, BoostDream can refine it into a high-quality 3D asset efficiently. In each set of images, the image on the left is the coarse 3D asset generated by Shap-E Jun and Nichol (2023a), and the three images on the right are our refined 3D asset.
@@ 1 Introduction
The significance of 3D assets has grown remarkably due to the expansion of virtual reality and the gaming industry. Creating these assets, however, involves considerable labor costs. Recently, the popularity of methods based on differentiable rendering Mildenhall et al. (2021); Shen et al. (2021); Wang et al. (2021); Kerbl et al. (2023) and the explosion of text-to-image models Rombach et al. (2022); Zhang et al. (2023a) provide a new line of methods for generating 3D assets.
Plenty of research Nichol et al. (2022); Jun and Nichol (2023a); Poole et al. (2022); Wang et al. (2023b); Lin et al. (2023); Chen et al. (2023b) in the text-to-3D domain aim to lower 3D asset creation costs, demonstrating the efficacy of leveraging 2D generation models for 3D tasks. Currently, there are two primary strategies in the field of text-to-3D generation. The first, exemplified by Point-E and Shap-E Nichol et al. (2022); Jun and Nichol (2023a), employs a feed-forward approach, using 3D datasets or latent representation mapped by the 3D datasets to train diffusion models. This enables quick, text-prompt-based 3D asset generation. However, these methods grapple with the limited variety and scope of 3D datasets.
In response to the constraints posed by limited 3D datasets, the SDS-based optimization method was developed. Models like DreamFusion and Magic3D Poole et al. (2022); Lin et al. (2023) exemplify this approach. They employ an SDS loss function Poole et al. (2022), enabling text-to-image diffusion models to train 3D assets predominantly represented by methods based on differentiable rendering, such as NeRF Mildenhall et al. (2021). This approach allows for the generation of high-quality 3D assets, leveraging the strong text comprehension and 2D generation abilities of diffusion models. However, this method is hindered by slow generation speeds, as each text prompt requires training a 3D representation from scratch and suffering from the Janus (multi-head) problem. How to efficiently generate high-quality 3D assets remains an urgent problem to be solved.
@@ 2.1 Feed-Forward Generation Method
Feed-forward generation method tries to directly utilize the diffusion models in text-to-3D tasks, which trained diffusion models with 3D assets as datasets Liu et al. (2023c, a); Sanghi et al. (2023); Nichol et al. (2022); Zhou et al. (2021) or perform diffusion with a latent representation of 3D assets Gupta et al. (2023); Ntavelis et al. (2023); Chen et al. (2023a); Zeng et al. (2022). Shap-E Jun and Nichol (2023a) is a typical example of this type of approach. It trains a 3D encoder to transform 3D assets to a set of parameters of an implicit function and then trains a diffusion model on these parameters. Compared to the methods using frozen 2D diffusion models, these approaches can generate 3D assets at a very fast speed due to their feed-forward nature. However, these methods were trained with the 3D datasets, while the 3D dataset available today is relatively small and low-quality. Consider the dataset used in 2D image generation tasks, Laion5B Schuhmann et al. (2022) contains more than 5 billion image-text pairs while the largest 3D dataset available, Objaverse-XL Deitke et al. (2023) can only carry 10 million 3D assets with worse quality captions. This causes the 3D generation results to be less sharp, have lower fidelity, and be unable to produce assets with complex semantics.
@@ 2.2 SDS-Based Optimization Generation Method
Inspired by the great success of using diffusion models in text-to-image tasks, one mainstream of works focuses on using 2D diffusion models along with differentiable 3D representing techniques like NeRF Mildenhall et al. (2021) to facilitate 3D generation. Dreamfields Jain et al. (2022) used pre-trained CLIP models to optimize NeRF, hoping to get a visually real result. But it takes much time and computing resources to generate. DreamFusion Poole et al. (2022) proposes SDS loss to optimize NeRF representation. SJC Wang et al. (2023a) is a concurrent effort of Dreamfusion, with the difference that SJC applies the chain rule to the 2D score. Research on providing supervision of novel views via SDS loss is becoming a trend Lin et al. (2023); Chen et al. (2023b); Metzer et al. (2023); Liu et al. (2023b); Qian et al. (2023); Melas-Kyriazi et al. (2023); Lorraine et al. (2023). Among them, ProlificDreamer Wang et al. (2023b) models the 3D parameters as variable, proposes variable fraction distillation (VSD) to address the shortcomings in SDS. These works indicate the capability to produce high-quality outputs enriched with detailed semantic information derived from 2D diffusion models. However, these SDS-based methods are impeded by prolonged optimization periods and the occurrence of Janus problems.
@@ 2.3 Multi-View in 3D Generation
The multi-view aware diffusion method in image generation has made great progress recently. MVDiffusion Tang et al. (2023) is capable of generating consistent multi-view images from text prompts given pixel-to-pixel correspondences with global awareness. The success of such a method also draws attention to 3D generation tasks. EfficientDreamer Zhao et al. (2023) used a 2D diffusion model that can generate an image consisting of four orthogonal-view sub-images and use it in the 3D generation. It has been proven to be efficient in improving the quality of 3D content and alleviating the Janus problem. MVDream Shi et al. (2023) further discusses the situation of building a multi-view diffusion model from both 2D and 3D data. It can benefit from the generalizability of the 2D diffusion model while keeping the consistency of 3D rendering, solving the Janus problem and content drift problem across different views. However, after training by the 3D datasets, the realistic texture information that the 2D diffusion model contained can be degraded. In our work, a multi-view rendering method is designed and a multi-view SDS loss function is formulated to guide the differentiable 3D representation, mitigating the Janus problem while obtaining high-quality generated results.
@@ 3.1 Background
For text-to-image tasks, Ho et al. (2020) has proposed a simplified training goal of Denoising Diffusion Probabilistic Models (DDPM). It is expressed as:
@@
In text-to-3D tasks using text-prompt as the generation condition, DreamFusion Poole et al. (2022); Wang et al. (2023b) has proposed the SDS loss. For a differentiable 3D rendering represented as $g(\theta,c)$, where $\theta$ are its parameters and $c$ are the camera parameters. Let $q_{t}^{\theta}\left(\mathbf{x}_{t}\mid c\right)$ be the distribution from diffusing $\mathbf{x}_{0}=g(\theta,c)$ to moment $t$. The optimization goal of SDS is by optimizing $\theta$, make the marginal distribution of the diffusion process of $q_{t}^{\theta}\left(\mathbf{x}_{t}\mid c\right)$, approach the marginal distribution of a pre-trained text-to-image model $p\left(\mathbf{x}_{t}\mid y\right)$ where $y$ is text-prompt. The training goal of the SDS loss method can be represented as:
@@
The idea of adding other control conditions in BoostDream is inspired by ControlNet Zhang et al. (2023b), which is a text-to-image task solution allowing more input as task specific conditions like images. Its training goal can be written as:
@@ 3.3 Multi-View Render System
$\mathbf{R}(\mathbf{a},\alpha)=\mathbf{I}+(\sin\alpha)\mathbf{K}+(1-\cos\alpha)\mathbf{K}^{2}$ (6)
@@
We use separate renderers for each task in different 3D representations. For a renderer $g$, we use the above four camera positions $p_{i}$ along with other parameters to get camera parameters $c_{i}$. With a total of four sub-images generated from different camera parameters $c_{i}$ and parameters of a 3D representation $\theta$, stitch these four images together to create one large $2\times 2$ composite image ${G}$. We also capture the normal maps rendered from 3D representation and follow a similar routine to get a large $2\times 2$ normal map ${N}$.
@@ 3.4 Multi-View SDS
Here, $\epsilon_{\phi}$ denotes the noise predicted by the diffusion model, $\lambda\in[0,1]$ balances the normal map conditioned and unconditioned noise predictions Chen et al. (2023c), and $s$ is the scale of classifier-free guidance Ho and Salimans (2022). $N$ represents the multi-view normal maps captured following the rule we proposed in section 3.3, for normal map renderer $g_{N}$, we use the same camera parameters used by composite image ${G}$ in Eq. 9:
@@
Furthermore, We enhance the loss with orientation loss $\mathcal{L}_{\mathrm{orient}}$ and opacity loss $\mathcal{L}_{\mathrm{opacity}}$ proposed at Poole et al. (2022), when implement with NeRF. The final loss function combines these components to ensure accurate surface detail representation, correct surface orientations, and optimal opacity:
@@
Figure 3: The first column is the Shap-E Jun and Nichol (2023a) results and the remaining column is the refined results of our method. The results show that BoostDream can refine and edit 3D assets according to different prompts based on input 3D assets.
@@ 4.1 Implementation Details
All the experiments in this paper are conducted on a single NVIDIA V100 GPU with 32GB VRAM. We use Shap-E Jun and Nichol (2023a) from official implementation Jun and Nichol (2023b) to generate coarse 3D assets and fit them into differentiable 3D representations in the first 200 iterations. After initialization, we set 4800 iterations for the refining process, of which the first 1800 iterations are under the guidance of the coarse 3D assets, and the subsequent 3000 iterations are guided by the differentiable rendering result itself. We use ControlNet 1.1 with Stable Diffusion 1.5 Zhang et al. (2023b) as the diffusion model. In the comparison experiment, we carried out experiments on different differentiable 3D representations, including NeRF Mildenhall et al. (2021), DMTet Shen et al. (2021), and 3D Gaussian Splatting Kerbl et al. (2023). Please refer to the Appendix Yonghao et al. (2024) for more details and results.
Figure 4: Comparision with Shap-E Jun and Nichol (2023a), DreamFusion Poole et al. (2022) and Magic3D Lin et al. (2023) for the same text-to-3D generation task. Our model has significantly stronger prompt relevancy and much better quality (best viewed by zooming in). See the results of our method on DMTet Shen et al. (2021) and 3D Gaussian Splatting Kerbl et al. (2023) in the Appendix Yonghao et al. (2024)
@@ 4.3 Comparison Experiment
Our BoostDream is a refining method that combines the advantages of the feed-forward generation method and the SDS-based optimization generation method. To demonstrate the effectiveness of our methods, we selected methods from each of the above two categories for the comparison experiment. For the feed-forward generation method, we choose Shap-E, which has an official implementation Jun and Nichol (2023b). The 3D assets generated by Shap-E are also used to initialize our BoostDream method. For the SDS-based optimization generation method, we choose DreamFusion and Magic3D, which are implemented by threestudio Guo et al. (2023) because no official code is available now. It is also worth noticing there are some differences between the original papers and the threestudio implementation. In this implementation, DreamFusion and Magcic3D use DeepFloyd StabilityAI (2023) as the diffusion model, while the originals use Imagen Saharia et al. (2022) and eDiff-I Balaji et al. (2022). We use the default parameter of these methods. The experiment result is shown in Figure 4. Our method can generate higher-quality 3D assets compared to the above methods. In rows 1, 4, and 7, DreamFusion and Magic3D have arisen with the Janus problem, which means the Gundam has multiple arms, the horse has wrong legs, and the elephant has numerous heads, while our method can alleviate this problem due to the three-stage refine.
@@ 4.4 Ablation Study
We also conducted two other ablation studies. One uses the coarse 3D asset generated by the feed-forward method to initialize the traditional SDS-based method. Another is using different control conditions to replace normal maps, such as canny edge and depth maps. See more ablation studies in the Appendix Yonghao et al. (2024).

### v4

@@ Abstract
Figure 1: Comparison of 3D Generation Results of baseline and BoostDream. Provided with a coarse 3D asset and text prompt pair, BoostDream can refine it into a high-quality 3D asset efficiently. In each set of images, the image on the left is the coarse 3D asset generated by Shap-E jun2023shape, and the three images on the right are our refined 3D asset.
@@ 1 Introduction
The significance of 3D assets has grown remarkably due to the expansion of virtual reality and the gaming industry. Creating these assets, however, involves considerable labor costs. Recently, the popularity of methods based on differentiable rendering mildenhall2021nerf; shen2021dmtet; wang2021neus; kerbl3Dgaussians and the explosion of text-to-image models rombach2022high; zhang2023adding provide a new line of methods for generating 3D assets.
Plenty of research nichol2022pointe; jun2023shape; poole2022dreamfusion; wang2023prolificdreamer; lin2023magic3d; chen2023fantasia3d in the text-to-3D domain aim to lower 3D asset creation costs, demonstrating the efficacy of leveraging 2D generation models for 3D tasks. Currently, there are two primary strategies in the field of text-to-3D generation. The first, exemplified by Point-E and Shap-E nichol2022pointe; jun2023shape, employs a feed-forward approach, using 3D datasets or latent representation mapped by the 3D datasets to train diffusion models. This enables quick, text-prompt-based 3D asset generation. However, these methods grapple with the limited variety and scope of 3D datasets.
In response to the constraints posed by limited 3D datasets, the SDS-based optimization method was developed. Models like DreamFusion and Magic3D poole2022dreamfusion; lin2023magic3d exemplify this approach. They employ an SDS loss function poole2022dreamfusion, enabling text-to-image diffusion models to train 3D assets predominantly represented by methods based on differentiable rendering, such as NeRF mildenhall2021nerf. This approach allows for the generation of high-quality 3D assets, leveraging the strong text comprehension and 2D generation abilities of diffusion models. However, this method is hindered by slow generation speeds, as each text prompt requires training a 3D representation from scratch and suffering from the Janus (multi-head) problem. How to efficiently generate high-quality 3D assets remains an urgent problem to be solved.
@@ 2.1 Feed-Forward Generation Method
Feed-forward generation method tries to directly utilize the diffusion models in text-to-3D tasks, which trained diffusion models with 3D assets as datasets liu2023meshdiffusion; liu2023one2345; sanghi2023clip; nichol2022pointe; zhou20213d or perform diffusion with a latent representation of 3D assets gupta20233dgen; ntavelis2023autodecoding; chen2023single; zeng2022lion. Shap-E jun2023shape is a typical example of this type of approach. It trains a 3D encoder to transform 3D assets to a set of parameters of an implicit function and then trains a diffusion model on these parameters. Compared to the methods using frozen 2D diffusion models, these approaches can generate 3D assets at a very fast speed due to their feed-forward nature. However, these methods were trained with the 3D datasets, while the 3D dataset available today is relatively small and low-quality. Consider the dataset used in 2D image generation tasks, Laion5B schuhmann2022laion5b contains more than 5 billion image-text pairs while the largest 3D dataset available, Objaverse-XL deitke2023objaverse can only carry 10 million 3D assets with worse quality captions. This causes the 3D generation results to be less sharp, have lower fidelity, and be unable to produce assets with complex semantics.
@@ 2.2 SDS-Based Optimization Generation Method
Inspired by the great success of using diffusion models in text-to-image tasks, one mainstream of works focuses on using 2D diffusion models along with differentiable 3D representing techniques like NeRF mildenhall2021nerf to facilitate 3D generation. Dreamfields jain2022dreamfields used pre-trained CLIP models to optimize NeRF, hoping to get a visually real result. But it takes much time and computing resources to generate. DreamFusion poole2022dreamfusion proposes SDS loss to optimize NeRF representation. SJC wang2023score is a concurrent effort of Dreamfusion, with the difference that SJC applies the chain rule to the 2D score. Research on providing supervision of novel views via SDS loss is becoming a trend lin2023magic3d; chen2023fantasia3d; metzer2023latent; liu2023zero; qian2023magic123; melas2023realfusion; lorraine2023att3d. Among them, ProlificDreamer wang2023prolificdreamer models the 3D parameters as variable, proposes variable fraction distillation (VSD) to address the shortcomings in SDS. These works indicate the capability to produce high-quality outputs enriched with detailed semantic information derived from 2D diffusion models. However, these SDS-based methods are impeded by prolonged optimization periods and the occurrence of Janus problems.
@@ 2.3 Multi-View in 3D Generation
The multi-view aware diffusion method in image generation has made great progress recently. MVDiffusion tang2023mvdiffusion is capable of generating consistent multi-view images from text prompts given pixel-to-pixel correspondences with global awareness. The success of such a method also draws attention to 3D generation tasks. EfficientDreamer zhao2023efficientdreamer used a 2D diffusion model that can generate an image consisting of four orthogonal-view sub-images and use it in the 3D generation. It has been proven to be efficient in improving the quality of 3D content and alleviating the Janus problem. MVDream shi2023mvdream further discusses the situation of building a multi-view diffusion model from both 2D and 3D data. It can benefit from the generalizability of the 2D diffusion model while keeping the consistency of 3D rendering, solving the Janus problem and content drift problem across different views. However, after training by the 3D datasets, the realistic texture information that the 2D diffusion model contained can be degraded. In our work, a multi-view rendering method is designed and a multi-view SDS loss function is formulated to guide the differentiable 3D representation, mitigating the Janus problem while obtaining high-quality generated results.
@@ 3.1 Background
For text-to-image tasks, NEURIPS2020_4c5bcfec has proposed a simplified training goal of Denoising Diffusion Probabilistic Models (DDPM). It is expressed as:
@@
In text-to-3D tasks using text-prompt as the generation condition, DreamFusion poole2022dreamfusion; wang2023prolificdreamer has proposed the SDS loss. For a differentiable 3D rendering represented as $g(\theta,c)$, where $\theta$ are its parameters and $c$ are the camera parameters. Let $q_{t}^{\theta}\left(\mathbf{x}_{t}\mid c\right)$ be the distribution from diffusing $\mathbf{x}_{0}=g(\theta,c)$ to moment $t$. The optimization goal of SDS is by optimizing $\theta$, make the marginal distribution of the diffusion process of $q_{t}^{\theta}\left(\mathbf{x}_{t}\mid c\right)$, approach the marginal distribution of a pre-trained text-to-image model $p\left(\mathbf{x}_{t}\mid y\right)$ where $y$ is text-prompt. The training goal of the SDS loss method can be represented as:
@@
The idea of adding other control conditions in BoostDream is inspired by ControlNet controlnetmodel, which is a text-to-image task solution allowing more input as task specific conditions like images. Its training goal can be written as:
@@ 3.3 Multi-View Render System
$\mathbf{R}(\mathbf{a},\alpha)=\mathbf{I}+(\sin\alpha)\mathbf{K}+(1-\cos\alpha)\mathbf{K}^{2}\$(6)
@@
We use separate renderers for each task in different 3D representations. For a renderer $g$, we use the above four camera positions $p_{i}$ along with other parameters to get camera parameters $c_{i}$. With a total of four sub-images generated from different camera parameters $c_{i}$ and parameters of a 3D representation $\theta$, stitch these four images together to create one large $2\times 2$ composite image G{G}. We also capture the normal maps rendered from 3D representation and follow a similar routine to get a large $2\times 2$ normal map N{N}.
@@ 3.4 Multi-View SDS
Here, $\epsilon_{\phi}$ denotes the noise predicted by the diffusion model, $\lambda\in[0,1]$ balances the normal map conditioned and unconditioned noise predictions chen2023control3d, and $s$ is the scale of classifier-free guidance ho2022classifierfree. $N$ represents the multi-view normal maps captured following the rule we proposed in section 3.3, for normal map renderer $g_{N}$, we use the same camera parameters used by composite image G{G} in Eq. 9:
@@
Furthermore, We enhance the loss with orientation loss $\mathcal{L}_{\mathrm{orient}}$ and opacity loss $\mathcal{L}_{\mathrm{opacity}}$ proposed at poole2022dreamfusion, when implement with NeRF. The final loss function combines these components to ensure accurate surface detail representation, correct surface orientations, and optimal opacity:
@@
Figure 3: The first column is the Shap-E jun2023shape results and the remaining column is the refined results of our method. The results show that BoostDream can refine and edit 3D assets according to different prompts based on input 3D assets.
@@ 4.1 Implementation Details
All the experiments in this paper are conducted on a single NVIDIA V100 GPU with 32GB VRAM. We use Shap-E jun2023shape from official implementation ShapE2023 to generate coarse 3D assets and fit them into differentiable 3D representations in the first 200 iterations. After initialization, we set 4800 iterations for the refining process, of which the first 1800 iterations are under the guidance of the coarse 3D assets, and the subsequent 3000 iterations are guided by the differentiable rendering result itself. We use ControlNet 1.1 with Stable Diffusion 1.5 controlnetmodel as the diffusion model. In the comparison experiment, we carried out experiments on different differentiable 3D representations, including NeRF mildenhall2021nerf, DMTet shen2021dmtet, and 3D Gaussian Splatting kerbl3Dgaussians. Please refer to the Appendix Boost2024 for more details and results.
Figure 4: Comparision with Shap-E jun2023shape, DreamFusion poole2022dreamfusion and Magic3D lin2023magic3d for the same text-to-3D generation task. Our model has significantly stronger prompt relevancy and much better quality (best viewed by zooming in). See the results of our method on DMTet shen2021dmtet and 3D Gaussian Splatting kerbl3Dgaussians in the Appendix Boost2024
@@ 4.3 Comparison Experiment
Our BoostDream is a refining method that combines the advantages of the feed-forward generation method and the SDS-based optimization generation method. To demonstrate the effectiveness of our methods, we selected methods from each of the above two categories for the comparison experiment. For the feed-forward generation method, we choose Shap-E, which has an official implementation ShapE2023. The 3D assets generated by Shap-E are also used to initialize our BoostDream method. For the SDS-based optimization generation method, we choose DreamFusion and Magic3D, which are implemented by threestudio threestudio2023 because no official code is available now. It is also worth noticing there are some differences between the original papers and the threestudio implementation. In this implementation, DreamFusion and Magcic3D use DeepFloyd Deepfloyd as the diffusion model, while the originals use Imagen saharia2022Imagen and eDiff-I balaji2022ediffi. We use the default parameter of these methods. The experiment result is shown in Figure 4. Our method can generate higher-quality 3D assets compared to the above methods. In rows 1, 4, and 7, DreamFusion and Magic3D have arisen with the Janus problem, which means the Gundam has multiple arms, the horse has wrong legs, and the elephant has numerous heads, while our method can alleviate this problem due to the three-stage refine.
@@ 4.4 Ablation Study
We also conducted two other ablation studies. One uses the coarse 3D asset generated by the feed-forward method to initialize the traditional SDS-based method. Another is using different control conditions to replace normal maps, such as canny edge and depth maps. See more ablation studies in the Appendix Boost2024.
@@ 5 Conclusion
Appendix
Appendix A Implementation Details
In the training process of the NeRF mildenhall2021nerf setting, we set rendering resolution to $128\times 128$, and batch size to 1. We apply our random multi-view render system to capture a combined image with four sub-images with rotation angle $\alpha$ set to $90^{\circ}$. We use AdamW optimizer kingma2014adam with learning rate $1\times 10^{-2}$ and $1\times 10^{-3}$ for geometry and background modeling. The background is replaced with random colors with $80\%$ of chance. In the DMTet shen2021dmtet setting, most of the parameters stay the same, but in the self-boost stage, we increase the resolution to $512\times 512$ for a better result. The initialization stage of 3D Gaussian Splatting kerbl3Dgaussians is somehow different from the other two methods as they use hash-grid while 3D Gaussian Splatting is able to initialize from point cloud representation directly. The rendering resolution is also $512\times 512$.
We apply the CFG trick and negative prompts following the example from MVDream shi2023mvdream, further append prompt “, 3d asset” or “, multi-view of the 3d asset” to get a more consistent result.
Figure 7: Simply Combination Ablation Study. The first column is the input coarse model generated by Shap-E jun2023shape, while the next three columns are the results for the original DreamFusion poole2022dreamfusion, the DreamFusion with initialization stage, and our BoostDream-NeRF, respectively.
Appendix B Simply Combination Ablation Study
Our BoostDream method does not just simply combine the feed-forward approach with the SDS-based method. To further test the benefits of applying our multi-view based strategy, we also design an ablation study using DreamFusion poole2022dreamfusion with the same initialization stage as our method. We use the results from Shap-E jun2023shape in the initialization stage and use the same prompt text as input to optimize the NeRF representation with DeepFloyd Deepfloyd. The results of the original DreamFusion, the DreamFusion with initialization stage, and our BoostDream-NeRF are shown in Figure 7. We can see in the first row even with the proper initialization, DreamFusion still suffers from the Janus problem and has coarse results compared to our BoostDream results.
Appendix C Control Condition Ablation Study
We also test our method with different multi-view control conditions replacing the normal map. We choose canny edge canny1986computational and depth map ranftl2020towards as guidance obtained through the same multi-view render system as normal map. The results are shown in the Figure 8. Canny edge just contains the edge information of the 3D asset. Intuitively, it is unsuitable as a control condition when generating high-quality 3D assets. The results also illustrate this point: when using canny edge as the control condition, the 3D asset suffers from incomplete generation. Especially in the second row, the bear turns out to be unnatural and has strange colors. Instead of canny edge using edge information to guide the refinement process, the depth map utilizes depth information, leading to complete generation results. However, we find that the generated results are less detailed when the control condition is depth map. This can be explained by the fact that minor details information is not prominent in depth map but salient in normal map zhang2023adding. We can further validate this idea with the last column, the generated 3D assets are high-quality and with more details when under the guidance of normal map.
Figure 8: Control Condition Ablation Study. The first column is the input coarse model generated by Shap-E jun2023shape, while all other columns are the output of our BoostDream method with different control conditions.
Appendix D Result on Different 3D Representations
This section supplements the comparison experiment in Section 4.3. We implement our BoostDream on other differentiable representations, including DMTet shen2021dmtet and 3D Gaussian Splatting kerbl3Dgaussians. The results are shown in Figure 9, illustrating the generality of our method in generating high-quality assets using different differential 3D representations.
Figure 9: Result on Different 3D Representations.The first column is the input coarse model generated by Shap-E jun2023shape, while the next three columns are the results of our BoostDream method implemented with NeRF mildenhall2021nerf, DMTet shen2021dmtet and 3D Gaussian Splatting kerbl3Dgaussians, respectively.